OpenStudy (asadkarim7):

the tangent and normal to the curve y= sin x at the point P where x= π/3 cut the x-axis at A and B respectively (a). show that AB = (5√3)/4. (b). find the area of the triangle PAB.

OpenStudy (amistre64):

the tangent to the curve is what..cos(pi/3) right?

OpenStudy (amistre64):

the normal is 90 from that....

OpenStudy (amistre64):

i dont quite understand the cuts the x axis at A and B

OpenStudy (anonymous):

maybe x and y axis at A and B?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

they mean the tangent cuts at a, and the normal cuts at b

OpenStudy (asadkarim7):

let me post the digaram

OpenStudy (anonymous):

its very simple find the eqns of the lines , find the intercepts of those lines , then use distance formula

OpenStudy (amistre64):

lol.... yeah, if you post it that would be more helpful :)

OpenStudy (amistre64):

simple yes, once you understand the question ;)

OpenStudy (anonymous):

just distance formula and area of a triangle = (1/2) bh , thats all there is , you dont even need a picture for this, should be able to draw one yourself, or visualise in your mind pretty easily

OpenStudy (amistre64):

well the tangent is simply its derivative; cos(pi/3)

OpenStudy (amistre64):

the equation of the line then becomes: y-(sin(pi/3)) = cos(pi/3)(x-(pi/3))

OpenStudy (amistre64):

cos(pi/3) = 1/2 right?

OpenStudy (amistre64):

the normal will have a slope of -2 then

OpenStudy (amistre64):

y-(sin(pi/3)) = -2(x-(pi/3)) for the normal equation

OpenStudy (asadkarim7):

Guys i still cant understand it , can sun1 explain it to me step by step ? plz

OpenStudy (amistre64):

they give you a point (pi/3, sin(pi/3)); or rather (pi/3, sqrt(3)/2) right?

OpenStudy (anonymous):

first take the derivative of sine, which is cosine. then plug in \[\frac{\pi}{3}\] into the derivative to get the slope. \[cos(\frac{\pi}{3})=\frac{1}{2}\] so that is the slope of the tangent line

OpenStudy (amistre64):

at this level you should have some redimentary skills at a graph id assume :)

OpenStudy (amistre64):

use the poit-slope form of a line to get the euation for the tangent and the norml

OpenStudy (anonymous):

amistre is in right direction,find the eqn of tangent and normal to the curve, find the x intercepts of both to find out the coordiantes of the two points, point P is given, it forms the triangle ABP

OpenStudy (amistre64):

someones been rearragnging my keyboard lol

OpenStudy (anonymous):

the equation of the line use the point slope formula. the point is \[(\frac{\pi}{3},\frac{\sqrt{3}}{2})\] and the slope is \[\frac{1}{2}\] so the equation is \[y-\frac{\sqrt{3}}{2}=\frac{1}{2}(x-\frac{\pi}{3})\]

OpenStudy (anonymous):

this crosses the x axis where \[y=0\] so set \[y=0\] and solve to get \[x=\frac{\pi}{3}-\sqrt{3}\] if i did the algebra right

OpenStudy (anonymous):

repeat the process for the line with slope -2 through \[(\frac{\pi}{3},\frac{\sqrt{3}}{2})\]

OpenStudy (amistre64):

x/2 - pi/6 = -sqrt.3/2 x = -sqrt3 + pi/3 its good ;)

OpenStudy (anonymous):

amistre did i screw this up? it looks good to me

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

the normal just uses -2 for th eslope then

OpenStudy (anonymous):

\[y-\frac{\sqrt{3}}{2}=-2(x-\frac{\pi}{3})\] is the normal line yes?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

let \[y=0\] get \[x=\sqrt{3}+\frac{\pi}{3}\]

OpenStudy (anonymous):

at least that is what i got

OpenStudy (amistre64):

/ -2 to begin with to ease the process

OpenStudy (amistre64):

sqrt.3/4 + pi/3 i think

OpenStudy (anonymous):

\[-\frac{\sqrt{3}}{2}=-2(x-\frac{\pi}{3})\] \[\sqrt{3}=x-\frac{\pi}{3}\] \[x=\sqrt{3}+\frac{\pi}{3}\]

OpenStudy (anonymous):

whoooooooooooooooooops

OpenStudy (anonymous):

lorda mercy

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{4}+\frac{\pi}{3}\]

OpenStudy (anonymous):

another dumb guy award for me.

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

lol.... thats what all my medals are for ;)

OpenStudy (anonymous):

ok fine. now that we have the x-intercepts, after some elementary algebra mistakes on my part, what are we supposed to do with them?

OpenStudy (amistre64):

throw tham at moving trains :)

OpenStudy (anonymous):

u both deserve for it :)

OpenStudy (anonymous):

i forgot what we were supposed to do with these. multiply? add? find the length?

OpenStudy (anonymous):

i mean for medals ;)..not throwing at the moving train :P

OpenStudy (amistre64):

the distance between them is just large - small

OpenStudy (anonymous):

now the final task, area of the triangle

OpenStudy (amistre64):

since they are on the number line; distance between them is large - small

OpenStudy (amistre64):

distance between 6 and -3 is 6--3 = 9 ;)

OpenStudy (amistre64):

height = sqrt3/2 ; distance = l-s --------------------------- = area 2

OpenStudy (anonymous):

meant to find an area lol but you guys have done it all over the place and its highly likely that the OP has just giving up on trying to follow what you guys have wrote

OpenStudy (amistre64):

the journey taken is more important the the final destination....... grasshopper

OpenStudy (amistre64):

sats been consistently and systematically finding the points... its good

OpenStudy (anonymous):

ooh some arithmetic i can do! \[\frac{\sqrt{3}}{4}+\frac{\pi}{3}-(\frac{\pi}{3}-\sqrt{3})=\frac{\sqrt{3}}{4}+\sqrt{3}=\frac{5\sqrt{3}}{4}\]

OpenStudy (anonymous):

wow just what it said it was!

OpenStudy (anonymous):

very east to set out working out to make the solution of these questions look really easy instead of making it look so complicated with 1000s of posts etc.

OpenStudy (anonymous):

true but very hard to type all in in one box. specially if typesetting.

OpenStudy (amistre64):

depends on if you wanna read a novel or have a conversation type environment; there isnt only one way to do things :)

OpenStudy (anonymous):

hi yo man

OpenStudy (amistre64):

this type of environment is suited best for interaction between questioneer and answeer

OpenStudy (amistre64):

if they really want a static read thru answer; theres always the internet lol

OpenStudy (anonymous):

Asadkarim7 i will be happy to work all this out on one paper and send it to you if this is confusing. the idea is this: find the slope of the tangent line, then find its equation. find the slope of the normal line, and its equation. then in each equation set y = 0 and solve for x. that is the whole idea. once you have both x intercepts subtract to get the distance. the rest is mechanics, but i will be happy to write it out in all its gory details without my algebra errors if you like

OpenStudy (amistre64):

asad is intelligent :)

OpenStudy (anonymous):

i am sure he is but i will be happy to do it to practice my latex

OpenStudy (anonymous):

make a pdf

OpenStudy (amistre64):

:) latex.... i perfer ascii lol

OpenStudy (asadkarim7):

thanks :P amistre64

OpenStudy (anonymous):

luddite

OpenStudy (anonymous):

^ "its very simple find the eqns of the lines , find the intercepts of those lines , then use distance formula" esessentially what I wrote like a whole hour ago

OpenStudy (amistre64):

ohhh..... so everything after that is useless and should be erased becasue its just utter garbage? .... comes across a bit pompuos dontcha think ;)

OpenStudy (amistre64):

in fact the "i still dont understand can you explain it to me step by step" kinda leads me to believe otherwise ... lol

OpenStudy (anonymous):

also, geez, way to spam the question guys, nw my computer is going extremely slow when I view this particular question because you guys spammed it so much :|

OpenStudy (amistre64):

youre prolly using internet explorer; i had to down firefox to avoid that

OpenStudy (amistre64):

the equation editor stuff tends to put a strain on the processor; which is why i tend to avoid it

OpenStudy (amistre64):

so asad; how can we help you with this problem?

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

let me know if you have any questions

OpenStudy (amistre64):

you forgot to put in the area lol

OpenStudy (anonymous):

oh heavens. fine.

OpenStudy (amistre64):

and here my fine useless garbage contribution ;)

OpenStudy (asadkarim7):

sattelite73 u made the equation y = sqrt 3/ 2 it should have been y- sqrt 3/2

OpenStudy (anonymous):

\[\frac{1}{2}bh=\frac{1}{2}\times \frac{5\sqrt{3}}{4}\times \frac{\sqrt{3}}{2}=\frac{15}{16}\]

OpenStudy (amistre64):

its his latex.... makes him go a little insane at times and starts putting = signs for - signs :)

OpenStudy (anonymous):

everyone is a critic. ok that was a typo and look, it is gone!

OpenStudy (amistre64):

that ones just blank.....

OpenStudy (asadkarim7):

u r genius sattelite 73

OpenStudy (anonymous):

what lovely font! what nice equations!

OpenStudy (anonymous):

how didi u make pdf file

OpenStudy (anonymous):

eventually i will learn this latex ... eventually.

OpenStudy (amistre64):

he starts out with some fine kindling, then adds a bit of tinder to feed the flame..

OpenStudy (anonymous):

i am using texmaker as an editor. just converts for you if you ask, and then can ask for external viewer. really quite friendly. i tried to use this on micordemon windows whatever and it took me maybe a day and a half to download what i needed. miktex and all that garbage. never worked right. left the darkside for linux a few months ago, to maybe 2 minutes do download and works like a charm.